OpenStudy (askme12345):

a statistics exam had test scores with a normal distribution. the mean score was a 75 with a standard deviation of 4

OpenStudy (askme12345):

what is the probability that a randomly selected student has a test score between 65 and 90

OpenStudy (anonymous):

familiar of reading z-score table?

OpenStudy (askme12345):

yes have one in front of me now

OpenStudy (askme12345):

not sure how to use it with this question though

OpenStudy (anonymous):

do you know how to find the z-score?

OpenStudy (askme12345):

i know it is sample value-sample mean/ standard deviation but i dont know what to enter for the sample value

OpenStudy (anonymous):

ok, what are the two samples values in this case?

OpenStudy (askme12345):

65 and 90

OpenStudy (askme12345):

do i do the average of the two

OpenStudy (anonymous):

find the z-score for each sample value

OpenStudy (askme12345):

-2.5 for 65 and 3.75 for 90

OpenStudy (askme12345):

so now do i find those numbers on the z chart? for -2.5 i would have .0062 and for 3.75 i would have .9999

OpenStudy (anonymous):

so P = P(z=3.75) - P(z=-2.5) will give you the answer

OpenStudy (askme12345):

so .9999-.0062 gives me .9937

OpenStudy (askme12345):

the next question asks what test score gives you the 95th percentile i would do the revese of this problem correct

OpenStudy (anonymous):

yea

OpenStudy (askme12345):

can u walk me through one more?